Tech Staff Meeting Notes February 5, 2008

Current Top Priorities are: Station maintenance and GPS wrapup (e.g. 110 foot tower at NMKM) followed closely by TDOT wireless link.

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

The task list updated February 6, 2008. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


Station Maintenance, NMSZ

Station maintenance is the top priority.

Holly's list on 2/5/2008:
TIPT (all)
EPRM (all)
CHRM (all)


From 120 seconds at 050505

Note that QUAR traces were actually JHAR data for practically all of 2007. Jim will get precise dates to Holly. Holly will then find and nuke any bad QUAR trace data as well as QUAR picks.

Vandalism update. GRAT is back . Need to look for a spot on NWCC property to move TIPT (see the Warden).

Need a dual mux assembly or two. Have the AMUX64T's. Pat will build up the rest of the hardware.
MKTA spare air conditioner is repaired. Needs to be returned to the doghouse.
Need a remote reset at lnxt. We'll use an ADAM 6060 internet reset device.
Need new RF preamps, radio align, etc, etc at MKTA.
PEBM and HENM sometimes have telemetry fades. Should reevaluate comms.
Brush at NMAD isn't too bad, we may just need to trim up around the guy anchors.

Mud Island vertical array is operational.

Steve and Chris attended the TDOT planning meeting with BoB Parrish yesterday. Steve supplied these notes:

Just to let everyone know the results of the planning meeting that Chris Mac, and I attended yesterday with Bob Parish. The main result was that it will likely be mid May before things will be ready for us to install our equipment at the Boswell site. The contractor will provide a route for the two coaxial cables coming into the the TMC hub building (near the comm tower). They will also provide optical fiber from the Hub building to the TMC building. The TMC building will contain the video wall and the T1 channel bank and router card. We will need to provide an ethernet2fiber media converter for both the Hub and TMC ends of the fiber. The fibers will be "62.5 multi-mode". Hope that is meaningful to someone.

Bob Parish was interested in someone from CERI attending these planning meetings on a regular basis (bi-weekly?) This could be useful if there is the possibility that we might make additional use of TDOT's network infrastructure. It also sounded like they might be amenable to our using additional band-width in the future.

We want to get as much of this engineered and completed as possible before the end of March (Steiner retires beginning of April).

The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos?
Lost contact with recorders 1 and 3. Can reset all 7 recorders which implies the equipment enclosure at each node has power and that at least the reset radio and telemetry are working.


We're in the process of getting James out of the station maintenance business. GPS stations will be treated as any of the seismic network stations. Mac will be the lead for routine maintenance. Data handling and processing is still done by Bob et al.

Greg is in charge of getting anything done to the GPS stations to Steinerize them. This is necessary to make sure we can maintain them without burning alot of time on reapplying bandaids. Note that they are in pretty good shape already. The primary hassle is communications.

RLAP, LCHS, and NWCC will be telemetered to NMKM. There will be a 110 foot tower at NMKM that links to the DBB at NMAD. Plan B is using a satellite based ISP ad NMKM. Plan C is sending everything to the water tower at Tiptonville and get on the DBB from there.

Items left:
Comms for MAIR, CJTR.
Portageville needs a visit to check router enclosure for fan and general SOH. Also need to replace 6A power supply with a 3A.


Notes from John

To be installed:

LRVA - We have permission, can install any time.

WSNC - (Winespring Bald, NC) - Special use permit turned into Forest Service, waiting for their O.K.

Maryville - After performing a radio test, it was found that the cave site we had permission for will not make the radio link to Starr Mtn. We can either set up a repeat or find another site. Given the other problems with the cave site, (possible noise problem from swimming pool pump, extremely low clearance, would have to modify an S-13 to get it to fit), I think it is worth looking for another location. I have gotten several possible cave sites from cavers in the area.


CCRT - Needs a rebuild after a bad lightning strike. I have the Panda II box here and am realigning it.

PLVA - Martin Chapman wants to take it over for VPI since we have abandoned it, I suspect though he won't be able to make the radio shot because we are on the south side of the mountain.

CBSC - USC wants to take this station over as a regional station. The USGS seems to have dropped it from their list of backbone stations.

NC Node - Need to figure out how to fix GPS time drop-outs. Greg suggests moving the antenna away from the building. I will go in the spring, the Forest Service doesn't recommend I try to drive up there unless I really need to, apparantly the road is passable but not recommended.

GFM - Need a crew to go back up and repair storm damage and replace presumably bad pre-amp card.

RBNC - Replace battery(s), and add another solar panel. Same deal as NC Node concerning roads though. I tried to go up last week and even though I had a key, I had to stop at the first tunnel because of ice.

CPCT - Re-work lighting and pest chaser.

WMTN - Should we consider digging it up and doing another pit (even though I swore I wouldn't do that) The existing one is eroding away and I just don't know how long it's going to last. We may have to end up having this one in sediment. This is the one Mitch and Greg helped dig the pit for, it's so steep on the side of that hill, that the top side of the pit is 10 ft. deep and the bottom side is 6 ft.

ETT Node - Passed USFS inspection, but TN forest service wants us to do work around their shop in the valley? Need to rebuild/repair remote reset. We are still not sure if the problem is at Starr Mtn. or at the EMA office. Need to decide whether it would be better to build something in the shop and take it up or try to repair on site.

GMG - E-W is bad, need to check out. Need battery box solar reg mounting plate and need to fuse the battery.

SWET - Chris found the freewave enclosure full of water, we need to design a rain shield for all the aluminum Hoffman boxes, they are leaking at the hinged side because the weather-stripping isn't making good contact with the frame. He put some good tape on it. The battery box was in good shape.

BHT - I have never heard anything from the FAA about moving the station to their site on Black Mtn. to improve access and get stuff out of a tree. Right now everything is in a new tree and working great, so there is no rush until that tree dies and falls over.

MVKY - Wait until construction is done and assist Bob Wurth in repairing/rebuilding it.

NCCU - Greg has the CMG-6 and is working on it.

C1SC and C2SC - The lantronix boxes are in non air conditioned rooms and routinely (especially C2SC) fail in the summer. I've found that if you pull the mss 100 out, wait for 1/2 hour and put them back in, they work fine until it heats up again. Can I modify a box with a cooling fan? Is there another device that deals with the heat of Charleston in the summer?

MYNC - NSN station that has never worked well and doesn't work at all now. USGS is debating on what to do with it. Need to contact them and ask them to make a decision because we are relying on it to plug a hole in our network. If they abandon it, what will we do? Short period at BBG?

Network Completion

The map shows stations with yellow triangles being new ones.
SMNC4done (see note above)
BHT8move (eventually)

Routine Maintenance Kudzu, battery check, etc, etc.

Need to swap out the old node computers at NHIN, GNAR, and AZTN. Steve needs to go to NHIN (but Steve isn't allowed to do anything else until the ew 7.2 upgrade is complete).

Second Tier Priorities

CVTN optical isolation. Need to upgrade to year 3 hardware. Hardware ready.

Steve spoke with Dean Kemp at NWS. They are willing to let us piggyback on their cable internet (not Dean is looking into the possibility of a static IP. This would get NAIT online. Bill called him Dean and didn't have answers to some questions. He'll call him back armed with the IT fact sheet for the mss100's.

Lenox Agreement. Mitch spoke with Ron Martens, EMF Broadcasting. There is a desire for some sort of MOU and a visit to CERI. Mitch will set this up.

FDL. This is the FDL Policy. .
Numerous cables from the vandalism need repaired.
Need 2 firewire cables. They are $500 each from the manufacturer and generic ones don't appear to work. The speed isn't sufficient. So what's special about the expensive ones? Bill will try to find out.

As Time Permits

3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. Chris will check it out to assess the added lightning hazard and see if anything can be done about it. Might be hard to do without the wiring diagram. Chris will make one if needed as time permits.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Inventory Installing new smeagol broke mysql. Sid is helping get the tables hooked back into mysqld.
Jim has all the broadband sensors defined and is working on additional equipment (e.g. isis) and the connections. Mitch will install Bill's account in mysql and inv so he can start building stations. Mitch needs to sit down with Jim and figure out what we need to have ISTI help us with. Sid will be available after about mid-May to finish up anything we need his help to take care of. Apparently the downloaded docs from guralp for the cmg5td's have the old +-4g contants. Mitch has paper documentation with the correct +-2g constants.

Steiner Brain Dump Seminar Need more of them as weather permits. Especially calibration and computer room power, general station power, and shake table operation.


Need new doors on the USGS shed. A small child with a plastic screwdrive could get in. This topic fueled a discussion on security because Chris caught two guys red handed stealing his trailer. The got away without the trailer. Chris is in discussion with UoM Police for and additional call box. Several ideas were given to improve security but it seems the most cost effective is to make sure things are locked, and restricted from open view where possible.

Withers needs to determine current orientation for MPHB so we can pull it and orient correctly.

Need to teach someone (our hazmat dude?) to neutralize etchant and clean the etching tanks. Along those same lines, Greg has a ton of drawings that play only with the "Eagle" software on his computer. Mitch is unaware of a way to convert to a more generic format. Also require a printer that does negative (the one on Greg's serial port does, others do not). Mitch knew this dependence on vendor specific software would bite us one day. That day is soon.

Pink Palace hardware needs to be rebuilt. New construction on the lot on the northeast corner of Central and Highland will likely block the analog transmission to the Pink Palace. Its probably time for a digital display anyway (or perhaps an IRIS AS-1 seismometer). Mitch will ask Gary about it.

John needs a spectrum analyzer.

Evaluate computer room power. Steve will be the keeper of the spreadsheet post-Steiner.

Build more battery charge controllers.